OpenStudy (anonymous):

how do you solve x^4+13x^2=-36?

OpenStudy (mertsj):

Factor it.

OpenStudy (mertsj):

\[(x^2+9)(x^2+4)=0\]

OpenStudy (mertsj):

Now set each factor equal to zero and solve.

OpenStudy (mertsj):

\[x^2=-9\] \[x=\pm 3i\] \[x^2=-4\]

OpenStudy (mertsj):

\[x=\pm2i\]

OpenStudy (anonymous):

is x=+-2i the answer?

OpenStudy (mertsj):

Did you read what I wrote? There are 4 answers.

OpenStudy (precal):

x to the fourth power means you have 4 solutions